# Complements of discriminants of real parabolic function singularities. II

### v1

@@ 2.3. Components for the $X_{9}^{2}$ singularities
The group $G_{0}$ (see § 2.1) acts on the space $\Phi_{2}$ of all polynomials (1) with A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the complement of the discriminant in some way. All components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
@@
acts on the space $\Phi_{2}$ (and on the entire space of polynomials of degree four). This operation always alters the real connected components of the discriminant complement, as well as the associated virtual components of type $\Phi_{2}$. Indeed, it maps any generic polynomial with value $I$ of the invariant Ind to a polynomial with value 3I-3-I. This operation is of order eight. All even iterations of it act as rotations by angles of $\frac{\pi k}{2}$ and thus belong to the group $G_{0}$. So, the entire set of connected components of the discriminant complement in the space $\Phi_{2}$ splits into the orbits of this operation. The cardinalities of all these orbits are divisors of 8.
@@
There are exactly 18 virtual components of class $\Phi_{2}$. The shapes of the zero-level sets of the polynomials representing nine of these virtual components are shown in Fig. 3 (see also Fig. 33 in [18]), and the polynomials representing the remaining nine virtual components are obtained from these by the operation ((18)).
@@
Theorem 18 was conjectured in [12], and the difference of all the connected components mentioned there was proven. It remains to prove that there are no additional components, that is, the virtual components listed in Proposition 17 are associated with no more real components than indicated by this theorem.
@@
Each polynomial of degree four in ${\mathbb{R}}^{2}$ having a critical point of class $D_{6}$ with zero critical value and a critical point of class $A_{3}$ can be reduced by a linear transformation to the form
(19) $\pm\left[t\left(\tilde{x}^{2}\tilde{y}+\frac{1}{3}\tilde{x}^{3}\right)+\frac{3}{4}\tilde{x}^{4}+3\tilde{x}^{3}\tilde{y}+\frac{3}{2}\tilde{x}^{2}\tilde{y}^{2}-\tilde{x}\tilde{y}^{3}\right]\ ,\hskip 23.49976ptt>0\ .$
@@
Proof. Of all the polynomials (19), only this one has critical value $\frac{-1}{324}$. $\Box$
@@
If a polynomial $f(x,y)$ of degree four has critical points of classes $D_{6}$ and $A_{3}$ and principal homogeneous part
(21) $x^{4}+2Ax^{2}y^{2}+y^{4}\ ,$
then $A$ is a root of the equation
(22) $\hskip 136.5733pt\frac{(A^{2}+3)^{3}}{27(A^{2}-1)^{2}}=5.\hskip 136.5733pt\Box$
@@
whose zero-level set is topologically equivalent to the picture (d), but has a singular point of class $E_{6}$ at the point closest to the center of the picture. Then, we perturb this singular point by adding very small terms of lower degrees. These terms do not change the topology of the curve far away from this singular point, but provide the standard dissolutions of the $E_{6}$ singularity, see Figure 24 in [18]. All the resulting polynomials are invariant under the reflection in the axis $y=0$. Thus, in each case, the virtual component is associated with no more than four real connected components of the discriminant complement. These four components differ for the topological reasons, see [12].
For the remaining nine virtual components, the realizations of the corresponding real connected components that prove Theorem 18 in these cases can be obtained from the polynomials (23)–(31) via the operation (18). $\Box$
3. Components for $J_{10}$
@@
with ten parameters $\lambda_{1},\dots,\lambda_{9},$ and $A$, see [5].
Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (32) and (33), respectively.
@@
If the virtual components of two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are the same, then their connected components of the space of generic polynomials are either the same or mapped to each other by the involution
@@ 3.2. Components for $J_{10}^{1}$ singularities
There are exactly ten virtual components of the formal graph of $J_{10}^{1}$ singularities. These virtual components are represented by polynomials, the topological shapes of whose zero-level sets are shown on Fig. 4. $\Box$
@@
There are exactly thirteen connected components of the discriminant complement in the space of all polynomials of class $\Omega_{1}$. Each of the virtual components shown in pictures $(a)$, $(b)$, (b~)($\tilde${b}), $(c)$, (c~)($\tilde${c}), $(d)$, and (d~)($\tilde${d}) of Fig. 4 is associated with a single real connected component. Each of the virtual components/topological types shown in pictures $(e)$, (e~),($\tilde${e}), and $(f)$ is associated with two different components that are related by the involution ((34)).
Proof of Theorem 28. The topological types shown in pictures ($a$), ($b$), ($c$), and ($d$) are represented respectively by the following polynomials:
@@
The realizations of pictures (b~)($\tilde${b}), (c~)($\tilde${c}), and (d~)($\tilde${d}) can be obtained from the realizations of pictures $(b),$ $(c)$, and $(d)$ by the involution
@@
No polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(f)$ can belong to the same connected component as its image f(x,y)f(x,-y) under the involution (34). Indeed, the polynomials (32) are of degree three in the variable $x$. Therefore, any vertical line $\{y=\mbox{const}\}$ can intersect the curve $\{f=0\}$ at most three times.
An invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ realizing the picture $(e)$ from its mirror image f(x,y)f(x,-y) is as follows. Four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement. Let us take three arbitrary points inside the rightmost three ovals. There is a single parabola of the form $x=ay^{2}+by+c$ that passes through these three points. The restriction of the polynomial $f$ to this parabola is a polynomial of degree at most six in the coordinate $y$. This restriction has six roots, so it is of degree exactly six. This parabola cannot intersect the fourth oval, hence it passes either above or below it with respect to the $x$ coordinate. The position of the parabola relative to the fourth oval is independent of the choice of three points inside the ovals, and remains unchanged along the connected components of the discriminant complement. However, these relative positions are different for any polynomial with the topological shape of the zero-level set shown in picture $(e)$ and for its mirror image f(x,y)f(x,-y). The same is true for polynomials that realize the picture (e~)($\tilde${e}). $\Box$
@@ 3.3. Components for $J_{10}^{3}$ singularities
There are exactly 23 virtual components of class $\Omega_{3}$. The shapes of the zero-level sets of polynomials $f(x,y)$ representing some thirteen of them are shown in Fig. 5 (see also Fig. 29 in [18]). In all these cases except those shown in pictures $(h)$, $(l)$, and $(m)$, the polynomials $f(x,y)$ and f(x,y)-f(-x,y), which are related by the involution ((35)), represent different virtual components.
We will denote by (a~)($\tilde${a}), (b~)($\tilde${b}),…the pictures obtained from the corresponding pictures $(a)$, (b)(b),…by the involution induced by the transformation (35), that is, by the reflection in a horizontal line and simultaneous change of all signs.
@@
Pictures $(d)$ and $(e)$ of Fig. 5 are topologically equivalent, however it follows from the Bezout’s theorem that they represent different connected components.
@@
There are exactly 33 connected components of the complement of the discriminant variety of the versal deformation ((33)). Each virtual component represented by a picture of Fig. 5 in the list
@@
Proof. The difference of all these 33 connected components easily follows from the topological considerations and Bezout’s theorem, see [12]. By Proposition 25, it remains to be shown that, in all cases not covered by the list (36), the involution (34) does not change the real connected component.
@@
$(x-t)^{3}-(x-t)y^{4}+3(x-t)^{2}+5(x-t)y^{2}-4(x-t)-Ct^{3}$
with arbitrary $t>0$ realizes the picture $(a)$ for $C<6-\frac{14}{3}\sqrt{\frac{7}{3}}$, the picture $(b)$ for $C\in\left(6-\frac{14}{3}\sqrt{\frac{7}{3}},-\frac{1}{2}\right)$, and the picture $(d)$ for $C\in\left(-\frac{1}{2},0\right)$. All of these polynomials are even with respect to the variable $y$, so their virtual components are represented by single connected components of the complements of the discriminants.
Picture $(m)$ is represented by the polynomials
$x(x-y^{2}-t)(x+y^{2}+t),\qquad t>0,$
which are also self-symmetric.
@@
$(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+3t(x-t)y^{2}+Ct^{3}$
realize picture $(c)$ for any $C\in(0,1/2))$. They are also invariant under the involution (34).
The polynomials
$(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+5t(x-t)y^{2}-4t^{2}(x-t)+\frac{t^{2}}{10}y^{2}-0.35t^{3}$

### v2

@@ 2.6. Components for the $X_{9}^{2}$ singularities
The group $G_{0}$ (see § 2.3) acts on the space $\Phi_{2}$ of all polynomials (1) with $f_{A}=x^{4}+2Ax^{2}y^{2}+y^{4},$ A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the complement of the discriminant in some way. All these components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
@@
acts on the space $\Phi_{2}$. This operation always alters the real connected components of the discriminant complement, as well as the associated virtual components of type $\Phi_{2}$. Indeed, it maps any generic polynomial with value $i$ of the invariant Ind to a polynomial with value 3i-3-i. This operation is of order eight. All even iterations of it act as rotations by angles of $\frac{\pi k}{2}$ and thus belong to the group $G_{0}$. So, the entire set of connected components of the discriminant complement in the space $\Phi_{2}$ splits into the orbits of this operation.
@@
There are exactly 18 virtual components of class $\Phi_{2}$. The shapes of the zero-level sets of the polynomials representing nine of these virtual components are shown in Fig. 3 ((see also Fig. 33 in [24])), and the polynomials representing the remaining nine virtual components are obtained from these by the operation ((17)).
@@
Theorem 18 was conjectured in [18], and the difference of all the connected components mentioned there was proven. It remains to prove that there are no additional components, that is, the virtual components listed in Proposition 17 are associated with no more real components than indicated in this theorem.
@@
Each polynomial of degree four in ${\mathbb{R}}^{2}$ having a critical point of class $D_{6}$ with zero critical value and a critical point of class $A_{3}$ can be reduced by an affine transformation of ${\mathbb{R}}^{2}$ to the form
(18) $\pm\left(t\left(\tilde{x}^{2}\tilde{y}+\frac{1}{3}\tilde{x}^{3}\right)+\frac{3}{4}\tilde{x}^{4}+3\tilde{x}^{3}\tilde{y}+\frac{3}{2}\tilde{x}^{2}\tilde{y}^{2}-\tilde{x}\tilde{y}^{3}\right)\ ,\quad t>0\ .$
@@
Proof. Of all the polynomials (18), only the polynomial (19) has critical value $\frac{-1}{324}$. $\Box$
@@
If a polynomial $f(x,y)$ of degree four with principal homogeneous part
(20) $x^{4}+2Ax^{2}y^{2}+y^{4}$
has critical points of classes $D_{6}$ and $A_{3}$, then $A$ is a root of the equation
(21) $\frac{(A^{2}+3)^{3}}{27(A^{2}-1)^{2}}=5.$
Proof. This follows from the well-known formula for $j$-invariant of quartic forms, see e.g. [10].
@@
whose zero-level set is topologically equivalent to the picture (d), but has a singular point of class $E_{6}$ at its center. Then, we perturb this singular point by adding very small terms of lower degrees. These terms do not change the topology of the curve far away from this singular point, but provide the standard dissolutions of the $E_{6}$ singularity, see Figure 24 in [24]. All the resulting polynomials (2630) are invariant under the reflection in the axis $y=0$. Thus, in each of these five cases, the virtual component is associated with no more than four real connected components of the discriminant complement. These four components differ for the topological reasons.
For the remaining nine virtual components, the convenient realizations of the corresponding real connected components can be obtained from the polynomials (22)–(30) via the operation (17). $\Box$
3. Components for $J_{10}$ singularities
@@
with ten parameters $\lambda_{1},\dots,\lambda_{9},$ and $A$, cf. [10].
Definition 7.
A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of the form (31) or (32) is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, all of its critical values are different and not equal to 0, and the complex zero-level set of the principal quasihomogeneous part of $f$ (of degree six with weights $\deg x=2,\deg y=1$) consists of three different complex parabolas or two parabolas and one line in ${\mathbb{C}}^{2}$.
Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (31) and (32), respectively. The formal graph and its virtual components for $J_{10}$ singularities are defined in the same way as for the case $X_{9}$, except that the frames of vanishing cycles consists of 10 elements.
@@
If the virtual components of two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are the same, then their connected components of the space of generic polynomials either are the same or are mapped to each other by the involution
@@ 3.2. Components for $J_{10}^{1}$ singularities
There are exactly ten virtual components of the formal graph of type $J_{10}^{1}$. These virtual components are represented by polynomials, the topological shapes of whose zero-level sets are shown on Fig. 4. $\Box$
@@
There are exactly thirteen connected components of the discriminant complement in the space of all polynomials of class $\Omega_{1}$. Each of the virtual components shown in pictures $(a)$, $(b)$, (b~)($\tilde${b}), $(c)$, (c~)($\tilde${c}), $(d)$, and (d~)($\tilde${d}) of Fig. 4 is associated with a single real connected component. Each of the virtual components/topological types represented in pictures $(e)$, (e~),($\tilde${e}), and $(f)$ is associated with two different components that are related by the involution ((33)).
Proof. The topological types shown in pictures ($a$), ($b$), ($c$), and ($d$) are represented respectively by the following polynomials:
@@
The realizations of pictures (b~)($\tilde${b}), (c~)($\tilde${c}), and (d~)($\tilde${d}) can be obtained from these realizations of pictures $(b),$ $(c)$, and $(d)$ by the involution
@@
No polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(f)$ can belong to the same connected component as its image f(x,y)f(x,-y) under the involution (33). Indeed, the polynomials (31) are of degree three in the variable $x$, so any vertical line $\{y=\mbox{const}\}$ can intersect the curve $\{f=0\}$ at most three times. Therefore, the ovals placed on the different sides of the infinite component of the zero-level set cannot pass one above the other. On the other hand, the involution (34) moves the component of picture (f) to the same component as (13).
An invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ realizing the picture $(e)$ from its mirror image f(x,y)f(x,-y) is as follows. Four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement. Let us take three arbitrary points inside the rightmost three ovals. There is a single parabola (or line) of the form $x=ay^{2}+by+c$ that passes through these three points. The restriction of the polynomial $f$ to this parabola is a polynomial of degree at most six in the coordinate $y$. This restriction has six roots, so it is of degree exactly six. This parabola cannot intersect the fourth oval, hence it passes either above or below it with respect to the $x$ coordinate. The position of this parabola relative to the fourth oval is independent of the choice of three points inside the ovals, and remains unchanged along the connected components of the discriminant complement. However, these relative positions are different for any polynomial $f(x,y)$ with the topological shape of the zero-level set shown in picture $(e)$ and for its mirror image f(x,y)f(x,-y). The same is true for polynomials that realize the picture (e~)($\tilde${e}). $\Box$
@@ 3.3. Components for $J_{10}^{3}$ singularities
There are exactly 23 virtual components of the formal graph of type $J_{10}^{3}$. The shapes of the zero-level sets of some polynomials $f(x,y)$ representing thirteen of these components are shown in Fig. 5. In all these cases except those shown in pictures $(h)$, $(l)$, and $(m)$, the polynomials $f(x,y)$ and f(x,y)-f(-x,y), which are related by the involution ((34)), represent different virtual components. $\Box$
@@
Twelve topological types shown in these pictures are given in Fig. 29 of [24]. Pictures $(d)$ and $(e)$ of Fig. 5 are topologically equivalent, however they represent different virtual components.
We will denote by (a~)($\tilde${a}), (b~)($\tilde${b}),…the pictures obtained from the corresponding pictures $(a)$, (b)(b),… of Fig. 5 by the transformation (34), that is, by the reflection in a horizontal line and simultaneous change of all signs.
@@
There are exactly 33 connected components of the complement of the discriminant variety of the versal deformation ((32)). Each virtual component represented by one of ten pictures
@@
Proof. The difference of all these 33 connected components follows from the topological considerations and Bezout’s theorem, see [18]. By Proposition 25, it remains to be shown that, in all cases not covered by the list (35), the involution (33) does not change the real connected component. The following polynomials (36), (37), (38), (39), are all invariant under this involution, so their virtual components are represented by single connected components of the complements of the discriminants.
The polynomial
(36) $(x-t)^{3}-(x-t)y^{4}+3(x-t)^{2}+5(x-t)y^{2}-4(x-t)-Ct^{3}$
with arbitrary $t>0$ realizes the picture $(a)$ for $C<6-\frac{14}{3}\sqrt{\frac{7}{3}}$, the picture $(b)$ for $C\in\left(6-\frac{14}{3}\sqrt{\frac{7}{3}},-\frac{1}{2}\right)$, and the picture $(d)$ for $C\in\left(-\frac{1}{2},0\right)$.
Picture $(m)$ is represented by the polynomials
(37) $x(x-y^{2}-t)(x+y^{2}+t),\quad t>0.$
@@
(38) $(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+3t(x-t)y^{2}+Ct^{3}$
realize picture $(c)$ for any $C\in(0,1/2)$.
@@
(39) $(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+5t(x-t)y^{2}-4t^{2}(x-t)+\frac{t^{2}}{10}y^{2}-0.35t^{3}$